OpenStudy (anonymous):

Help me solve? I'm not even sure where to begin... x^2+5x+3=0

OpenStudy (anonymous):

\[x ^{2}+5x+3=0\]

OpenStudy (phi):

It looks like you need to use the quadratic formula http://www.purplemath.com/modules/quadform.htm or use "completing the square"

OpenStudy (phi):

Have you learned either of these ?

OpenStudy (anonymous):

Yes but I have not used them in a while...

OpenStudy (phi):

if you more problems like this, I would watch the video, because once you learn it, the problems go quickly.

OpenStudy (anonymous):

I have watched that video before.

OpenStudy (phi):

ok, can you say what is a, b, and c in x^2+5x+3=0

OpenStudy (phi):

if you don't know, watch the first 3 minutes again.

OpenStudy (anonymous):

\[a=x^2, b=5x, c=3\]

OpenStudy (phi):

close, but a , b and c are only the numbers, not the letters a=1 (the number in front of x^2 is 1 if you don't see a number) b=5 c=3

OpenStudy (anonymous):

\[1+5+3=0??\]

OpenStudy (anonymous):

Oh, you are a life saver! Thank you so much :)

OpenStudy (anonymous):

kk